OpenStudy (anonymous):

Help! 6-4(3-5)^2+30/5

OpenStudy (anonymous):

6 - 4(-2)^2 + 6 6-4(4) + 6 6-16+6 -4

OpenStudy (gorv):

6-4*(-2)^2+6 6-4*4+6=12-16=-4

OpenStudy (anonymous):

@easyaspi314 is right

OpenStudy (anonymous):

lol Thx

OpenStudy (anonymous):

welcome